please show all work/steps taken or any reasoning used in solving the problem.

thank you!
Answer #1

use lens equation

1/f = 1/i + 1/o

where i is image distance and o is object distance

(a) As image is virtual, i is negative

1/o = 1/f - 1/i

1/o = 1/12 + 1/17

o = 7.0344 cm

--------------------------------

(b) m = -i/o

m = -(-17) / 7.0344

m = 2.4166

-----------------------------

(c) height of object = height of image / 0.41666

height of object = 8 / 2.41666 = 3.31 mm

--------------------------------------

(d) image is upright (because it is virtual)

-----------------------------------

(e) object and image are on the same side.

--------------------------------------